OpenStudy (anonymous):

For what values of r does the function e^(rx) satisfy the equation y''+5y' -6y=0 (Note: r is a constant.) For what values of r does the function e^(rx) satisfy the equation y''+5y' -6y=0 (Note: r is a constant.) @Mathematics

myininaya (myininaya):

solve \[r^2+5r-6=0\]

myininaya (myininaya):

\[(r+6)(r-1)=0\]

myininaya (myininaya):

r=-6 r=1

OpenStudy (anonymous):

did you just take the derivative of e^rx with respect to r and the second derivative and then plug it in the equation?

myininaya (myininaya):

\[y=e^{-6x}+e^{1x}\] <- you can check this if you want

myininaya (myininaya):

\[ay''+by'+cy=0\] all you have to do to find value of r is solve \[ar^2+br+c=0\]

OpenStudy (anonymous):

okay i see thank you

myininaya (myininaya):

oh and to give a more general answer we say \[y=c_1 e^{-6x}+c_2e^{x}\] c_1,c_2 are constants